# Thermodynamic approach to quantum cooling limit of continuous Gaussian feedback

### v1

@@ Abstract
Introduction.— Measurement and feedback are fundamental tools that enable noise suppression, stabilization, optimal control, and adaptive control in diverse applications. In quantum systems, continuous measurement and feedback based on real-time data have been theoretically studied Wiseman and Milburn (2009); Jacobs and Steck (2006); Wiseman and Milburn (1993); Wiseman (1994); Doherty and Jacobs (1999); Bushev et al. (2006) and experimentally realized for qubit stabilization Vijay et al. (2012); Ristè et al. (2013), error correction Livingston et al. (2022), and cooling of oscillatory systems Rossi et al. (2018); Gieseler et al. (2012); Magrini et al. (2021); Kamba et al. (2023). In nanomechanical resonators Rossi et al. (2018) and optically trapped nanoparticles Gieseler et al. (2012); Magrini et al. (2021); Kamba et al. (2023); Gonzalez-Ballestero et al. (2021), state-based feedback control reduces vibrational energy, enabling cooling close to the ground state. In Gaussian systems, the quantum Kalman filter Belavkin (1995) provides optimal filtering that enhances the precision of feedback and has been demonstrated experimentally Magrini et al. (2021).
@@
In this Letter, we establish a fundamental bound of feedback cooling in Gaussian quantum systems. Specifically, we derive a second-law-like inequality that imposes a constraint on work extraction by a measure of quantum information flow called quantum-classical-transfer (QC-transfer) entropy Yada et al. (2022). Since the obtained bound is based on the kinetic temperatures of the system, it enables characterization of feedback cooling (i.e., reduction of the temperature of the system) and is achievable with a sufficiently large feedback gain along with using the quantum Kalman filter, in contrast to previously known bounds Yada et al. (2022). We numerically demonstrate these results with experimentally relevant parameters adopted from a recent nanoparticle experiment Magrini et al. (2021). Our results provide a theoretical foundation for the limits of feedback cooling from the perspective of thermodynamics of information in the quantum regime.
Figure 1: Schematic of the setup. The system undergoes feedback control based on estimated values obtained from the results of QND measurements. The system is also coupled to a heat bath at temperature $T$.
Setup.—We describe the framework for feedback cooling in Gaussian systems, where the state remains Gaussian throughout its evolution. We consider a quantum non-demolition (QND) continuous measurement and feedback control of a Gaussian system, which is also coupled to a single heat bath that satisfies the detailed balance condition (see Fig. 1). Specifically, the Hamiltonian of the system is supposed to be harmonic: $\hat{H}_{\mathrm{HO}}=\frac{\hbar\omega}{2}\left(\hat{p}^{2}+\hat{x}^{2}\right)$. Here, $\hat{x}$ and $\hat{p}$ represent the dimensionless position and momentum operators, respectively, satisfying the commutation relation $[\hat{x},\hat{p}]=i$, and $\omega$ is the angular frequency of the oscillator. We assume that the expectation values of the initial position and momentum are zero. The dynamics of feedback cooling are governed by the stochastic master equation Magrini et al. (2021):
@@
where $\mathcal{D}_{\hat{z}}[\hat{\rho}_{\mathrm{c}}]\coloneqq\hat{z}\hat{\rho}_{\mathrm{c}}\hat{z}^{\dagger}-\frac{1}{2}\{\hat{z}^{\dagger}\hat{z},\hat{\rho}_{\mathrm{c}}\}$ is the dissipator and $\hat{\rho}_{\mathrm{c}}(t)$ represents the conditional state at time $t$ given previous measurement outcomes. The superoperator $\mathcal{L}_{\mathrm{bath}}[\hat{\rho}_{\mathrm{c}}]\coloneqq\gamma(\bar{n}+1)\mathcal{D}_{\hat{a}}[\hat{\rho}_{\mathrm{c}}]+\gamma\bar{n}\mathcal{D}_{\hat{a}^{\dagger}}[\hat{\rho}_{\mathrm{c}}]$ represents the contribution of the heat bath, where $\hat{a}$ and $\hat{a}^{\dagger}$ represent the annihilation and creation operators, and $\gamma$ is the dissipation rate. The ratio between $\bar{n}$ and $\bar{n}+1$ satisfies the detailed balance condition $\bar{n}/(\bar{n}+1)=e^{-\hbar\omega/T}$, where $T$ is the temperature of the heat bath and the Boltzmann constant is set equal to unity for the sake of simplicity. The third term on the right-hand side of Eq. (Thermodynamic approach to quantum cooling limit of continuous Gaussian feedback) represents measurement backaction, while the fourth term reflects the reduction in position variance resulting from the measurement process. The superoperator $\mathcal{H}_{x}$ is defined as $\mathcal{H}_{\hat{x}}[\hat{\rho}_{\mathrm{c}}]\coloneqq\hat{x}\hat{\rho}_{\mathrm{c}}+\hat{\rho}_{\mathrm{c}}\hat{x}-2\mathrm{tr}(\hat{x}\hat{\rho}_{\mathrm{c}})\hat{\rho}_{\mathrm{c}}$. The parameter $k$ represents the measurement rate, while the Wiener increment $dW$ is a Gaussian random variable with zero mean and variance $dt$. The measurement result obtained via QND measurement of the position over a short time interval $dt$ is described by $dy(t)=\braket{\hat{x}(t)}_{\mathrm{c}}dt+dW/\sqrt{8k\eta}$, where $\eta$ represents the detection efficiency and $\braket{\cdot(t)}_{\rm{c}}\coloneqq\mathrm{tr}\left(\hat{\rho}_{\rm{c}}(t)\;\cdot\right)$ denotes the expectation value under the conditional state. The set of all measurement results until time $t$ is defined as $Y_{t}\coloneqq\{dy(s)\}_{s=0}^{t^{-}}$. Note that the unconditional state is given by $\hat{\rho}=\mathbb{E}_{Y_{t}}[\hat{\rho}_{\rm c}]$, where $\mathbb{E}_{X}$ represents taking the expectation value over a random variable $X$.
Estimates of the observable are constructed from the set of obtained measurement outcomes and used for feedback. Using measurement results $Y_{t}$, estimates are defined as $e_{t}=h(Y_{t})$, where $h(\cdot)$ compresses the past measurement data into estimates $e_{t}$. The posterior state conditioned on this estimate is denoted by $\hat{\rho}_{t}^{e_{t}}$. Using the estimates, the feedback action is determined. The linear feedback Hamiltonian $\hat{H}_{\rm{FB}}$ is given by
@@
QC-transfer entropy.— As our goal is to obtain an information-thermodynamic relation to characterize the cooling limit of Gaussian systems, we consider the QC-transfer entropy introduced in Ref. Yada et al. (2022), which quantifies the amount of information gain through continuous measurement. It is defined as
@@
where $\hat{\rho}_{\mathrm{c,\mathrm{M}}}:=\hat{\rho}_{\mathrm{c}}+2k\mathcal{D}_{\hat{x}}[\hat{\rho}_{\mathrm{c}}]dt+4k\eta\mathcal{H}_{\hat{x}}[\hat{\rho}_{\mathrm{c}}](dy(t)-\braket{\hat{x}(t)}_{\mathrm{c}}dt)$ is the conditional state after measurement at time $t$, and $S(\hat{\rho})\coloneqq-\mathrm{tr}[\hat{\rho}\ln\hat{\rho}]$ represents the von Neumann entropy. We note that Eq. (5) is regarded as the conditional QC-mutual information given all prior measurement results $Y_{t}$, because $\hat{\rho}_{\rm c}$ is the conditional state. Since the QC-transfer entropy quantifies the total information gain, it provides an upper bound for the information available in the multiple feedback processes.

### v2

@@ Abstract
I Introduction
Measurement and feedback are fundamental tools that enable noise suppression, stabilization, optimal control, and adaptive control in diverse applications. In quantum systems, continuous measurement and feedback based on real-time data have been theoretically studied Wiseman and Milburn (2009); Jacobs and Steck (2006); Wiseman and Milburn (1993); Wiseman (1994); Doherty and Jacobs (1999); Bushev et al. (2006) and experimentally realized for qubit stabilization Vijay et al. (2012); Ristè et al. (2013), error correction Livingston et al. (2022), and cooling of oscillatory systems Rossi et al. (2018); Gieseler et al. (2012); Magrini et al. (2021); Kamba et al. (2023). In nanomechanical resonators Rossi et al. (2018) and optically trapped nanoparticles Gieseler et al. (2012); Magrini et al. (2021); Kamba et al. (2023); Gonzalez-Ballestero et al. (2021), state-based feedback control reduces vibrational energy, enabling cooling close to the ground state. In Gaussian systems, the quantum Kalman filter Belavkin (1995) provides optimal filtering that enhances the precision of feedback and has been demonstrated experimentally Magrini et al. (2021).
@@ I Introduction
In this study, we consider a setup in which cooling is achieved by a continuous position measurement Jacobs and Steck (2006) together with feedback based on its measurement outcomes Wiseman and Milburn (2009); Doherty and Jacobs (1999) (see Fig. 1). In this framework, we establish a fundamental bound of feedback cooling in Gaussian quantum systems. Specifically, we derive a second-law-like inequality that imposes a constraint on work extraction by a measure of quantum information flow called quantum-classical-transfer (QC-transfer) entropy Yada et al. (2022). Since the obtained bound is based on the kinetic temperatures of the system, it enables characterization of feedback cooling (i.e., reduction of the temperature of the system) and is achievable with a sufficiently large feedback gain along with using the quantum Kalman filter, in contrast to previously known bounds Yada et al. (2022). We numerically demonstrate these results with experimentally relevant parameters adopted from a recent nanoparticle experiment Magrini et al. (2021). Our results provide a theoretical foundation for the limits of feedback cooling from the perspective of thermodynamics of information in the quantum regime.
This paper is organized as follows. In Sec. II, we introduce the setup for Gaussian feedback cooling and define key quantities including the transfer entropy and kinetic temperatures. In Sec. III, we present the main result, i.e., the information thermodynamic limit of quantum feedback cooling. In Sec. IV, we discuss the asymptotic achievement of the equality condition of the obtained bound. Finally, in Sec. V, we summarize our results and outline possible future directions.
Figure 1: Schematic of the setup. The system undergoes feedback control based on estimated values obtained from the outcomes of position measurements. The system is also coupled to a heat bath at temperature $T$.
II Setup
II.1 Model of continuous Gaussian feedback cooling
We describe the framework for feedback cooling in Gaussian systems, where the state remains Gaussian throughout its evolution. We consider continuous position measurement, whose measurement operator commutes with $\hat{x}$, and real-time feedback control based on the measurement record, applied to a Gaussian system coupled to a single heat bath that satisfies detailed balance (see Fig. 1). Specifically, the Hamiltonian of the system is supposed to be harmonic: $\hat{H}_{\mathrm{HO}}=\frac{\hbar\omega}{2}\left(\hat{p}^{2}+\hat{x}^{2}\right)$. Here, $\hat{x}$ and $\hat{p}$ represent the dimensionless position and momentum operators, respectively, satisfying the commutation relation $\hat{x},\hat{p}]=i$, and $\omega$ is the angular frequency of the oscillator. In addition, the system is externally controlled by a feedback Hamiltonian $\hat{H}_{\mathrm{FB}}(t)$ based on the past measurement outcomes. We assume that the expectation values of the initial position and momentum are zero. The dynamics of feedback cooling are governed by the stochastic master equation Jacobs and Steck (2006):
@@ II.1 Model of continuous Gaussian feedback cooling
where $\mathcal{D}_{\hat{z}}[\hat{\rho}_{\mathrm{c}}]\coloneqq\hat{z}\hat{\rho}_{\mathrm{c}}\hat{z}^{\dagger}-\frac{1}{2}\{\hat{z}^{\dagger}\hat{z},\hat{\rho}_{\mathrm{c}}\}$ is the dissipator and $\hat{\rho}_{\mathrm{c}}(t)$ represents a state conditioned on all measurement outcomes up to time $t$. The superoperator $\mathcal{L}_{\mathrm{bath}}[\hat{\rho}_{\mathrm{c}}]\coloneqq\gamma(\bar{n}+1)\mathcal{D}_{\hat{a}}[\hat{\rho}_{\mathrm{c}}]+\gamma\bar{n}\mathcal{D}_{\hat{a}^{\dagger}}[\hat{\rho}_{\mathrm{c}}]$ represents the contribution of the heat bath, where $\hat{a}$ and $\hat{a}^{\dagger}$ represent the annihilation and creation operators, and $\gamma$ is the dissipation rate. The ratio between $\bar{n}$ and $\bar{n}+1$ satisfies the detailed balance condition $\bar{n}/(\bar{n}+1)=e^{-\hbar\omega/T}$, where $T$ is the temperature of the heat bath and the Boltzmann constant is set equal to unity for the sake of simplicity. The third term on the right-hand side of Eq. (II.1) represents momentum diffusion at a rate $2k$, resulting from the backaction of the position measurement, while the fourth term represents the information obtained from the measurement process Jacobs and Steck (2006). Because the measurement operator $\sqrt{2k}\hat{x}$ commutes with $\hat{x}$, the measurement is a quantum-non-demolition (QND) position measurement if the Hamiltonian term $\hat{H}_{\mathrm{HO}}$ is absent. The superoperator $\mathcal{H}_{x}$ is defined as $\mathcal{H}_{\hat{x}}[\hat{\rho}_{\mathrm{c}}]\coloneqq\hat{x}\hat{\rho}_{\mathrm{c}}+\hat{\rho}_{\mathrm{c}}\hat{x}-2\mathrm{tr}(\hat{x}\hat{\rho}_{\mathrm{c}})\hat{\rho}_{\mathrm{c}}$. The parameter $k$ represents the measurement rate, while the Wiener increment $dW$ is a Gaussian random variable with zero mean and variance $dt$. The measurement outcome obtained via measurement of the position over an infinitesimal time interval $dt$ is described by $dy(t)=\braket{\hat{x}(t)}_{\mathrm{c}}dt+dW/\sqrt{8k\eta}$, where $\eta$ represents the detection efficiency and $\braket{\cdot(t)}_{\rm{c}}\coloneqq\mathrm{tr}\left(\hat{\rho}_{\rm{c}}(t)\;\cdot\right)$ denotes the expectation value under the conditional state. The set of all measurement outcomes until time $t$ is defined as $Y_{t}\coloneqq\{dy(s)\}_{s=0}^{t^{-}}$. Note that $\hat{\rho}_{\mathrm{c}}(t)$ is the state of the system for a given realization of measurement outcomes $Y_{t}$. We also denote $\hat{\rho}=\mathbb{E}_{Y_{t}}[\hat{\rho}_{\rm c}]$ as the averaged state by taking average over all measurement outcomes, where $\mathbb{E}_{X}$ represents taking the expectation value over a random variable $X$.
Estimates of the observable are constructed from the set of obtained measurement outcomes and used for feedback. Using measurement outcomes $Y_{t}$, estimates are defined as $e_{t}=h(Y_{t})$, where $h(\cdot)$ compresses the past measurement data into estimates $e_{t}$. The posterior state conditioned on this estimate is denoted by $\hat{\rho}_{t}^{e_{t}}$. Using the estimates, the feedback action is determined. The linear feedback Hamiltonian $\hat{H}_{\rm{FB}}$ is given by
@@
Note that obtaining the measurement outcomes reduces the uncertainty in position, although $\braket{\hat{x}}_{\mathrm{c}}$ and $\braket{\hat{p}}_{\mathrm{c}}$ fluctuates due to random measurement outcomes. The feedback Hamiltonian is then designed to decrease the absolute values of $\braket{\hat{x}}_{\mathrm{c}}$ and $\braket{\hat{p}}_{\mathrm{c}}$, and thus the fluctuation of position and momentum is reduced at the ensemble average level. Therefore, the dynamics described by Eq. (II.1) allows reducing the energy of the system and realizes feedback cooling.
II.2 QC-transfer entropy
As our goal is to obtain an information-thermodynamic relation to characterize the cooling limit of Gaussian systems, we consider the QC-transfer entropy introduced in Ref. Yada et al. (2022), which quantifies the amount of information gain through continuous measurement. It is defined as
@@ II.2 QC-transfer entropy
where $\hat{\rho}_{\mathrm{c,\mathrm{M}}}:=\hat{\rho}_{\mathrm{c}}+2k\mathcal{D}_{\hat{x}}[\hat{\rho}_{\mathrm{c}}]dt+4k\eta\mathcal{H}_{\hat{x}}[\hat{\rho}_{\mathrm{c}}](dy(t)-\braket{\hat{x}(t)}_{\mathrm{c}}dt)$ is the conditional state after measurement at time $t$, and $S(\hat{\rho})\coloneqq-\mathrm{tr}[\hat{\rho}\ln\hat{\rho}$] represents the von Neumann entropy. We note that Eq. (5) is regarded as the conditional QC-mutual information given all prior measurement outcomes $Y_{t}$, because $\hat{\rho}_{\rm c}$ is the conditional state. Since the QC-transfer entropy quantifies the total information gain, it provides an upper bound for the information available in the multiple feedback processes.
@@
II.3 Kinetic temperatures
We here introduce kinetic temperatures and work rate. We begin by defining the kinetic temperatures of a given Gaussian state. The one-mode Gaussian state is fully characterized by the mean values of position and momentum and the variance-covariance matrix. Let $\bm{\sigma}:=[\sigma_{ij}]_{i,j=x,p}$ denote the covariance matrix of the averaged state $\hat{\rho}(t)$. The position and momentum variances are given by $\sigma_{xx}\coloneqq\braket{\hat{x}^{2}}-\braket{\hat{x}}^{2}$ and $\sigma_{pp}\coloneqq\braket{\hat{p}^{2}}-\braket{\hat{p}}^{2}$, respectively, while the covariance is defined as $\sigma_{xp}\coloneqq\frac{1}{2}\braket{\{\hat{x},\hat{p}\}}-\braket{\hat{x}}\braket{\hat{p}}$, where $\langle\cdot\rangle$ denotes the expectation value with respect to $\hat{\rho}(t)$. Let the eigenvalues of $\bm{\sigma}$ be $\sigma_{uu}$ and $\sigma_{vv}$. Noting that a single-mode Gaussian state can always be expressed as a squeezed thermal state Kim et al. (1989), we define the kinetic temperatures along eigen-directions. The single-mode squeezed thermal state is generated by applying a squeezing operation $\hat{S}(\xi):=\exp\left(\frac{1}{2}\xi(\hat{a}^{\dagger})^{2}-\frac{1}{2}\xi^{*}\hat{a}^{2}\right)$ with a squeezing parameter $\xi=r\rm{e}^{i\theta}$ to the Gibbs state $\hat{\rho}^{\mathrm{G}}\coloneqq\exp(-\hat{H}_{\mathrm{HO}}/T^{\mathrm{G}})/\mathrm{tr}[\exp(-\hat{H}_{\mathrm{HO}}/T^{\mathrm{G}})]$ Olivares (2012), where $T^{\mathrm{G}}$ is the temperature of the Gibbs state, as $\hat{\rho}_{t}=\hat{S}(\xi)\hat{\rho}^{\mathrm{G}}\hat{S}(\xi)^{\dagger}$. The eigenvalues of $\bm{\sigma}$ satisfy $\sigma_{uu}={\rm{e}}^{2r}\sigma^{\rm{G}}$ and $\sigma_{vv}={\rm{e}}^{-2r}\sigma^{\rm{G}}$, where
$\sigma^{\mathrm{G}}:=\mathrm{tr}(\hat{x}^{2}\hat{\rho}^{\rm G})=\mathrm{tr}(\hat{p}^{2}\hat{\rho}^{\rm G})=\frac{1}{2}\coth\frac{\hbar\omega\beta^{\rm G}}{2},$ (6)
is the common position and momentum variance of the state $\hat{\rho}^{\mathrm{G}}$. Accordingly, the kinetic temperatures $T_{u}$ and $T_{v}$, associated with these eigenvalues, are defined as:
A schematic of kinetic temperatures is shown in Fig. 2. The kinetic temperatures $T_{u}$ and $T_{v}$ correspond to scaling the Gibbs state temperature
$T^{\mathrm{G}}=\hbar\omega\left[\ln\left(\frac{2\sqrt{\det\bm{\sigma}}+1}{2\sqrt{\det\bm{\sigma}}-1}\right)\right]^{-1},$ (8)
according to the widths in eigen-directions of the distribution. We further note that
$=-\frac{\hbar\omega\hat{u}^{2}}{2T_{u}}-\frac{\hbar\omega\hat{v}^{2}}{2T_{v}}-\ln Z(\beta^{\mathrm{G}}),$ (9)
where
$\hat{u}$ $\coloneqq e^{r}\hat{S}(\xi)\hat{x}\hat{S}(\xi)^{\dagger},$ (10)
$\hat{v}$ $\coloneqq e^{-r}\hat{S}(\xi)\hat{p}\hat{S}(\xi)^{\dagger},$ (11)
We next define the work rate as follows. First, the internal energy of the system is given by $E(t)=\mathrm{tr}[\hat{H}_{\mathrm{HO}}\hat{\rho}(t)]$. Then, based on the first law of thermodynamics, the work rate corresponding to the energy extracted from the system is defined as $\dot{W}_{\mathrm{ext}}:=\dot{Q}-\frac{d}{dt}E(t)$, where
$\dot{Q}:=\mathrm{tr}\left(\hat{H}_{\mathrm{HO}}\mathcal{L}_{\mathrm{bath}}[\rho]\right),$ (12)
is the heat Gong et al. (2016). The time variation of the energy can be expressed as contributions along the directions of $\sigma_{uu}$ and $\sigma_{vv}$, using the quadrature operators $\hat{u}$ and $\hat{v}$ associated with these directions, as $\frac{d}{dt}E(t)=\frac{\hbar\omega}{2}\frac{d}{dt}\mathrm{tr}\left[\hat{\rho}\hat{u}^{2}\right]+\frac{\hbar\omega}{2}\frac{d}{dt}\mathrm{tr}\left[\hat{\rho}\hat{v}^{2}\right]$. Similarly, the work rate can be further decomposed as:
$\dot{W}_{\mathrm{ext}}=\dot{W}_{\mathrm{ext},u}+\dot{W}_{\mathrm{ext},v},$ (13)
where the terms $\dot{W}_{\mathrm{ext},u}$ and $\dot{W}_{\mathrm{ext},v}$ represent the contributions of the work rate along the directions of $\sigma_{uu}$ and $\sigma_{vv}$, defined by Eqs. (50) and (51).
III Main results
III.1 Information thermodynamic cooling limit
We now present the main result of this study, a fundamental bound of the work rate in terms of kinetic temperatures. Suppose that the system is in the steady state; the averaged state $\hat{\rho}$ is given by the steady solution of Eq. (II.1). Then, we have a generalized second law of thermodynamics in the form
This inequality indicates how much QC-transfer entropy can be utilized to enhance work extraction and lower the kinetic temperatures, and serves as a quantum generalization of the second-law-like bound for classical Langevin systems Horowitz and Sandberg (2014). Since cooling is achieved by extracting work from the system and lowering its kinetic temperatures, inequality (14) represents the fundamental cooling limit.
We compare our result (14) with the generalized second law using the temperature $T$ of the heat bath, given in the form Yada et al. (2022):
$\frac{\dot{W}_{\rm{ext}}}{T}\leq\dot{I}_{\mathrm{QCT}}.$ (15)
In Sec. B.3 of the appendix, we show that the following relation holds in the steady state:
$\frac{\dot{W}_{\mathrm{ext}}}{T}\leq\frac{\dot{W}_{\mathrm{ext},u}}{T_{u}}+\frac{\dot{W}_{\mathrm{ext},v}}{T_{v}}.$ (16)
This relation shows that inequality (14) provides a tighter bound compared to (15) in the steady-state regime, since the former explicitly involves the kinetic temperatures of the system itself.
We next consider the conditions for achieving the equality in (14). A key ingredient is the quantum Kalman filter, which provides the conditional expectation values as the estimates, as presented in Eq. (3). Note that $\hat{\rho}_{t}^{e_{t}}=\hat{\rho}_{\mathrm{c}}$ holds in this case. By using this estimate based on the quantum Kalman filter, one can implement the so-called state-based feedback Doherty and Jacobs (1999); Magrini et al. (2021), where the feedback Hamiltonian (2) is given by the coefficients $c_{x}=b_{x}\langle\hat{x}\rangle_{\mathrm{c}}+b_{p}\langle\hat{p}\rangle_{\mathrm{c}}$ and $c_{p}=-a_{x}\langle\hat{x}\rangle_{\mathrm{c}}-a_{p}\langle\hat{p}\rangle_{\mathrm{c}}$. This feedback Hamiltonian aims to reduce the mean position and momentum to the origin in the phase space. Another requirement for achieving the equality in (14) is a large feedback gain. In fact, the equality in (14) is asymptotically achieved as the feedback gain goes to infinity if the estimate is implemented by the quantum Kalman filter (see Sec. IV for the proof). For simplicity, let the parameters of the feedback Hamiltonian be $a_{x}=b_{p}=g$ and $a_{p}=b_{x}=0$. In this case, the equality in (14) is asymptotically achieved in the limit $g\to\infty$. This implies that, under these conditions, the feedback system perfectly converts the acquired information into work.
Under the same conditions, the occupancy number of the harmonic Hamiltonian $\hat{H}_{\rm HO}$, $\langle a^{\dagger}a\rangle$, also reaches its minimal value determined by the bath temperature and the measurement rate (see Sec. IV), as the optimal feedback minimizes the sum of the variances of position and momentum, which is equivalent to the occupation number in our setup. This highlights the connection between the information-theoretic bound on thermodynamic energy conversion and the cooling limit of Gaussian systems.